OpenStudy (anonymous):

point on the curve 2y^3 + x^2 = 12y at which the tangent to the curve is vertical is

hartnn (hartnn):

vertical lines have slope =.... ?

hartnn (hartnn):

you get slope of tangent if you find dy/dx so first find it using implicit differentiation and equate it to slope of vertical line.

OpenStudy (anonymous):

can u plse solve it, i don;t know

hartnn (hartnn):

you don't know whats slope of vertical line ?? :O

hartnn (hartnn):

you know how to differentiate implicitly ?

OpenStudy (anonymous):

i don'tknow

OpenStudy (anonymous):

i know how to find dy/dx

OpenStudy (anonymous):

Slope of vertical line is either positive or negative infinity if I recall.

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{x}{6-3y^2}\]We seek the derivative to be infinity, so we have\[6=3y^2\]\[y=\pm\sqrt{2}\]Plugging the values of y into the original equation, we have\[2(\sqrt{2})^3+x^2=12\sqrt{2}\]and\[2(-\sqrt{2})^3+x^2=-12\sqrt{2}\]Solve for x for each corresponding value of y.

OpenStudy (anonymous):

Further calculation shows that \[12\sqrt{2}>2\sqrt{8}\] the first equation yield imaginary solution, which we don't really want. This implies only the second equation holds for real solution of x.

OpenStudy (anonymous):

Hope this helps